A sum of R8 000,00 is lent in the beginning of a year at a certain simple rate of interest. After eight months,


in a different loan, a sum of R1 164,00 more is lent but at the simple interest rate twice the former. At the


end of the year, R903,00 is earned as simple interest from both the loans. The original simple yearly rate of


interest, rounded to two decimal places, is

1
Expert's answer
2022-04-19T04:31:08-0400

Interest for 8 month = ("8000\cdot R \cdot 8)/(12 \cdot 100) =" 53R

Sum for last 4 month = "(8000+1164) = 9164"

Rate of the interest for last 4 month = 2R

Interest for 4 month = "(9164\cdot 2R \cdot 4)/(12 \cdot 100) = 61R"

Total interest for 1 year = 61R + 53R = 114R

114R = 903

R = 7.92

7,92%